40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Mar 9 10:57:06 2021, Paco, Anchal, Summary, IMC, XARM lock and POX spectra 20210309_POX11_Spec_XARMLocked.pdf20210309_XARM_Locked.tar.gz
    Reply  Tue Mar 9 14:37:26 2021, gautam, Summary, SUS, PRM suspension PRMtrip.png
       Reply  Tue Mar 9 15:22:56 2021, Koji, Summary, SUS, PRM suspension 
Message ID: 15887     Entry time: Tue Mar 9 14:37:26 2021     In reply to: 15884     Reply to this: 15889
Author: gautam 
Type: Summary 
Category: SUS 
Subject: PRM suspension 

The PRM got tripped ~5AM this morning. The cause is unclear - the seismometer reports elevated activity ~10 minutes before the ringdown starts (as judged using the OSEMs). But the other optics didn't seem to receive as much of an impulse (I only show the BS sensors here as it sits on the same stack as the PRM). Anyway it certainly wasn't me trying to make life difficult for the morning team.

I was able to restore the damping with reEnableWatchdogs.py. I am now running some suspension tests on the PRM by letting it swing freely so please let that finish. I plan to attempt some locking this evening.


[Paco, Anchal]

- Upon arrival, MC is locked, and we can see light in MON5 (PRM) (usually dark).

Attachment 1: PRMtrip.png  698 kB  | Hide | Hide all
ELOG V3.1.3-